Given the demand and supply equations,




Qdx = 150 - 20Px and Qsx=-30+50Px, Find a) Equilibrium price; b) if market price is Rs3.50, what will be amount of surplus units?



Plz i want a proper solution 🙏

1
Expert's answer
2021-12-19T18:09:22-0500

a) In equilibrium Qd = Qs, so:

"150 - 20Px = -30 + 50Px,"

70Px = 180,

Px = 2.57.

b) If market price is Rs 3.50, then the amount of surplus units is:

"Qs - Qd = (-30 + 50\u00d73.5) - (150 - 20\u00d73.5) = 65."